OpenStudy (kailee1423):

Find the volume of the square pyramid shown. Round to the nearest tenth if necessary. (Pic in comments)

OpenStudy (kailee1423):

@phi

OpenStudy (phi):

do you know the formula for the volume of a pyramid?

OpenStudy (phi):

It should be in your notes. or see http://www.basic-mathematics.com/volume-of-a-pyramid.html

OpenStudy (jhannybean):

Just curious, why is it divided by 3?

OpenStudy (kailee1423):

1/3Bh is the formula and I did it and got 55 but it was wrong

OpenStudy (phi):

B is short for the area of the base. in this case the area of a square with sides = 11

OpenStudy (kailee1423):

bh isnt base*height?

OpenStudy (jhannybean):

It is.

OpenStudy (kailee1423):

Oh there is something in my book about b being area of the base in a problem

OpenStudy (phi):

in the formula for the volume of a pyramid 1/3 Area of base times height the area of the base is sometimes abbreviated to B this is confusing. it does not mean the length of a side. It is just a "reminder" (which means you have to memorize this) B is AREA of the the base

OpenStudy (kailee1423):

So instead of B being 11, it would be 11*11?

OpenStudy (phi):

exactly

OpenStudy (kailee1423):

After doing that, the volume should be 605

OpenStudy (phi):

yes, 15/3 * 11*11 = 5*121= 605 putting in the units 605 cm^3 or 605 cc

OpenStudy (kailee1423):

Thanks again! I'll make another question lol

OpenStudy (phi):

**Just curious, why is it divided by 3?*** @Jhannybean I am sure we can derive it with calculus, and maybe just geometrically But I remember Archimedes found the volume of a cone was 1/3 the volume of the enclosing cylinder, and he was so proud of finding that, he had it engraved on his tombstone.

OpenStudy (jhannybean):

Oh, nice.